OpenStudy (anonymous):

Can you help me? Assign each letter and a blank space to a number as shown by the alphabet table below: 0 = _ 1 = A 2 = B 3 = C 4 = D 5 = E 6 = F 7 = G 8 = H 9 = I 10 = J 11 = K 12 = L 13 = M 14 = N 15 = O 16 = P 17 = Q 18 = R 19 = S 20 = T 21 = U 22 = V 23 = W 24 = X 25 = Y 26 = Z Use to encode the phrase “FIT AS A FIDDLE”.

OpenStudy (anonymous):



OpenStudy (anonymous):

What is the purpose of that matrix? I have no idea... The encoding is easy... just look up each letter and replace with the number. F = 6 I = 9 T = 20 (space) = 0 A = 1 S = 19 (space) = 0 A = 1 space = 0 F = 6 I = 9 D = 4 D = 4 L = 13 E = 5

OpenStudy (anonymous):

But again, is there more? I don't understand the matrix as part of this problem.

OpenStudy (anonymous):

what is all that you type up there ^^

OpenStudy (anonymous):

@malical

OpenStudy (anonymous):

Read down, from top to bottom... letters spell FIT AS A FIDDLE... numbers as the code numbers

OpenStudy (anonymous):

i dont have anything that looks like that

OpenStudy (anonymous):

and i really dont want to type all those numbers out how do i take a screen shot

OpenStudy (anonymous):

what format are your choices?

OpenStudy (anonymous):

are they lists of numbers?

OpenStudy (anonymous):

yes its a list of numbers

OpenStudy (anonymous):

screen shot is ALT and PRT SCR at the same time, usually.

OpenStudy (sirm3d):

i get the picture. the matrix will be used to encode "FIT AS A FIDDLE"

OpenStudy (anonymous):

@jake it didnt work:(

OpenStudy (anonymous):

Let's wait for these comments before continuing... maybe we'll understand the matrix.

OpenStudy (anonymous):

Moving onwards to screen shots? Alt + Prt Scr is Windows, Shift + Command + 4 is Mac. Does this have anything to do with the matrix you posted? Because that has negatives in it. Is that something you came up with or what the book told you would help? (Or the answer I guess.)

OpenStudy (anonymous):

its part of the question a. starts with 22 b. starts with 22 c. starts with 16 d. starts with -16

OpenStudy (anonymous):

A. 22 13 47 -12 3 13 -9 -19 0 5 29 -4 -2 -12 7

OpenStudy (anonymous):

B. 22 13 47 -12 3 13 -9 -19 0 -5 29 -4 -2 12 -7

OpenStudy (anonymous):

C. 16 5 17 -12 2 35 5 18 -12 8 61 6 32 -36 11

OpenStudy (anonymous):

D. -16 5 17 12 2 35 3 18 12 -8 61 6 -32 36 11

OpenStudy (anonymous):

you write the uncoded letters like I had above in a matrix Then you multiply it by the encoding matrix that the problem gives you, The solution is that matrix product.

OpenStudy (anonymous):

so the first row from my numbers is [ 6 9 20 ] (this represents "FIT") And you multiply that by the encoding matrix... maybe @sirm3d can help... I'm not good at that step.

OpenStudy (anonymous):

2nd unencoded row is "_AS" or [0 1 19] 3rd unencoded row is "_A_" or [0 1 0] 4th unencoded row is "FID" or [6 9 4] last row is "DLE" or [4 13 5]

OpenStudy (sirm3d):

like this \[\huge \left[\begin{matrix}2 & 1 & -1 \\ -1 & 1 & 0 \\0 & 2 & -1\end{matrix}\right]\left[\begin{matrix} 6 & 9 & 20 & 0 & 1 \\19 & 0 & 1 & 0 & 6 \\ 9 & 4 & 4 & 12 & 5\end{matrix}\right]\]

OpenStudy (anonymous):

oh, you do it as 3 x 5? I wasn't sure if it was that or 5 x 3.

OpenStudy (anonymous):

As I said, this isn't something I am good at :(

OpenStudy (sirm3d):

the first row should be 22 14 37 -12 3 not 22 13 47 -12 3

OpenStudy (anonymous):

oops, I've been calling "L" as 13, but it's 12 as @sirm3d shows it correctly

OpenStudy (sirm3d):

the 2nd and third rows are that of answer (A) so my money's on (A)